OpenStudy (anonymous):

A variable x is normally distributed with unknown mean and standard deviation =2. if the probability that x exceeds 7.5 is 0.8023, find the mean

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

im getting 9.2 am i correct??

Directrix (directrix):

My statistics is a bit rusty but I think the mean will not be greater than 7.5. z = (X - μ) / σ z = (7.5 - μ) / 2 The z score paired with .8023 is .805 which is above the mean. .805 = (7.5 - μ) / 2 μ = 5.89 Here's a link to the area under the curve chart: https://figures.boundless.com/18108/large/normal01.jpe